OpenStudy (anonymous):

y=3x-2 y=9x=4 Solve the system and show work please

OpenStudy (anonymous):

CAN SOMEBODY PLEASE HELP ME

OpenStudy (anonymous):

3x - 2 = 9x - 4 -2 + 4 = 9x - 3x -6 = 6x x = -1

OpenStudy (anonymous):

Thank You